what's the linear function f(4)=3 and f(6)=2
Answer by venugopalramana(3286)

what's the linear function f(4)=3 and f(6)=2
let the linear function in x be f(x)=ax+b
f(4)=4a+b=3....(1)
f(6)=6a+b=2....(2)
(2)-(1)....6a+b-4a-b=2-3=-1
2a=-1 or a=-1/2
substituting in (1)..we get 4*-1/2+b=3
-2+b=3 or b=3+2=5
hence the linear function is f(x)=ax+b = -x/2+5